OpenStudy (anonymous):

PLEASE HELP! Solve by completing the square: x^2-6x+8=0

OpenStudy (anonymous):

(x + a)(x+ b) = x^2 +(a+b)x +ab

OpenStudy (anonymous):

so (x-2)(x-4)

OpenStudy (solomonzelman):

\[x^2-6x+8=0\]\[x^2-6x+9=1\]\[(x-3)^{2}=1\]\[x-3=±1\] FINISH...

OpenStudy (anonymous):

wow lets try this start with \[x^2-6x=-8\] then since half of 6 is 3, and 3 squared is 9, to right to \[(x-3)^2=-8+9=1\] and go from there or the method of @SolomonZelman either way

OpenStudy (solomonzelman):

Well... what you said, @satellite73 is how my teacher taught me, but I only took what I did from that ;)

OpenStudy (wolf1728):

x² -6x +8=0 Move non-x term to the right x² -6x =-8 Divide the equation by the coefficient of X² The coefficient is 1 so this doesn't have to be done x² -6x =-8 take the coefficient of X which is -6 divide that by 2 which is -3 square that number 9 add it to both sides of the equation x² -6x +9= 1 Take square root of both sides x-3 = 1 x = 2

OpenStudy (wolf1728):

x = 4

OpenStudy (anonymous):

Thanks guys :) Much appreciated. Wish I could give more than one medal